OpenStudy (anonymous):

For which values of x are the given vectors u=xi−2j and v=xi+8j orthogonal?

OpenStudy (anonymous):

+ - 4

OpenStudy (anonymous):

The vectores are orthogonal if \[\langle u, v \rangle = x \cdot x + (-2) \cdot 8 = x^2 - 16 = 0\]so the solutions are\[x = \pm 4.\]

OpenStudy (anonymous):

i guess those weird brackets denote dot productions?

OpenStudy (anonymous):

Yes, those weird brackets denote dot product.